OpenStudy (anonymous):

5√6x+8√14x-9√6x+18√14x

OpenStudy (anonymous):

add like term \[5\sqrt{6}x-9\sqrt{6}x\]

OpenStudy (anonymous):

\[5\sqrt{6}x-9\sqrt{6}x+8\sqrt{14}x+18\sqrt{14}x\] \[26\sqrt{14}x-4\sqrt{6}x\]